The length of time a customer must wait in line at the post office has approximately a normal distribution with a mean of 5.8 minutes and a standard deviation of 2.6 minutes. Find the probability that a customer will wait between 4 and 8 minutes?
Answer by stanbon(75887)

The length of time a customer must wait in line at the post office has approximately a normal distribution with a mean of 5.8 minutes and a standard deviation of 2.6 minutes. Find the probability that a customer will wait between 4 and 8 minutes?
----
z(4) = (4-5.8)/2.6 = -0.6923
---
z(8) = (8-5.8)/2.6 = 0.8462
========================
P(4 < x < 8) = P(-0.6923 < z < 0.8462) = 0.7556
================
Cheers,
Stan H.
================